find the solution to the system using elimination
2x + 3y = 15
5x + 3y = 12

Answer by mathslover(157)

2x + 3y = 15
3y = 15 -2x (subtract 2x from both sides)
also given 5x + 3y = 12
Replacing 3y = 15 -2x in the equation above we get
5x + 15 -2x = 12
3x = -3 ( group the like terms at one side)
x = -1
3y = 15 - 2*(-1)
3y = 17
y=17/3
solution set
x=-1
y=17/3